# Robust First and Second-Order Differentiation for Regularized Optimal Transport

### v1

@@ 3 Differentiation of Loss Functions
$\frac{\partial\text{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{\partial\theta_{i}}$ $=\sum_{k=1}^{M}\frac{\partial{\boldsymbol{y}}_{k}}{\partial\theta_{i}}\frac{\partial\text{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{\partial{\boldsymbol{y}}_{k}},$ (13)
$\frac{\partial^{2}\mathrm{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{\partial\theta_{i}\partial\theta_{j}}$ $=\sum_{k=1}^{M}\sum_{s=1}^{M}\left(\frac{\partial{\boldsymbol{y}}_{s}}{\partial\theta_{i}}\right)\frac{\partial^{2}\mathrm{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{\partial{\boldsymbol{y}}_{s}\partial{\boldsymbol{y}}_{k}}\left(\frac{\partial{\boldsymbol{y}}_{k}}{\partial\theta_{j}}\right)^{\top}$
$\quad+\sum_{k=1}^{M}\frac{\partial^{2}{\boldsymbol{y}}_{k}}{\partial\theta_{i}\partial\theta_{j}}\frac{\partial\mathrm{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{\partial{\boldsymbol{y}}_{k}}.$ (14)
Here $C_{\theta}$ is the abbreviation of $C^{\boldsymbol{Y}_{\theta}\rightarrow\boldsymbol{Y}^{*}}$ emphasizing the dependence on $\boldsymbol{Y}_{\theta}$. The key step is to find the explicit expression for the first derivatives with respect to the source data $\frac{\partial\text{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{\partial{\boldsymbol{y}}_{k}}\in\mathbb{R}^{d}$ and the second-order derivatives with respect to source data $\frac{\partial^{2}\mathrm{OT}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})}{\partial{\boldsymbol{y}}_{s}\partial{\boldsymbol{y}}_{k}}\in\mathbb{R}^{d\times d}$.
@@ 3.1 Previous methods computing first- and second-order derivatives
For the analytical expression of the first-order derivatives of regularized OT distances, there are two main approaches: the Danskin approach and the implicit differentiation approach. The Danskin approach computes the gradient of EOT, based on applying the Danskin’s theorem to the dual function [5, 8, 15, 14]. Recall the Lagrangian $\mathcal{L}(C,\Pi,\boldsymbol{f},\boldsymbol{g})$ is a function of $C,\Pi,\boldsymbol{f}$ and $\boldsymbol{g}$ and $\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})=\max_{\Pi,\boldsymbol{f},\boldsymbol{g}}\mathcal{L}(C,\Pi,\boldsymbol{f},\boldsymbol{g})$. The Danskin’s theorem states that given $\boldsymbol{f}^{*},\boldsymbol{g}^{*},\boldsymbol{\Pi}^{*}$, the gradient of $\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ with respect to $\boldsymbol{y}_{k}$ is
$\frac{\partial\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{\partial\boldsymbol{y}_{k}}$ $=\sum_{ij}\frac{\partial C_{ij}}{\partial\boldsymbol{y}_{k}}\frac{\partial\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{\partial C_{ij}}=\sum_{ij}\frac{\partial C_{ij}}{\partial\boldsymbol{y}_{k}}\frac{\partial\mathcal{L}(C,\Pi^{*},\boldsymbol{f}^{*},\boldsymbol{g}^{*})}{\partial C_{ij}}$
$=\sum_{ij}\frac{\partial C_{ij}}{\partial\boldsymbol{y}_{k}}\Pi^{*}_{ij}=\sum_{j}\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\Pi^{*}_{kj}.$ (15)
@@
Once $\Pi^{*}$ is obtained from Sinkhorn algorithm, the first-order derivative with respect to the source dataset $\boldsymbol{Y}$ immediately follows without additional computational cost. The Danskin’s approach naturally extends to the de-biased Sinkhorn divergence $S_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})$, but unfortunately this approach doesn’t work on the Sinkhorn distances $\widetilde{\text{OT}}_{\epsilon}(C_{\theta},\boldsymbol{\mu},\boldsymbol{\nu})$ because it is not of the form $\max_{\Pi,\boldsymbol{f},\boldsymbol{g}}\phi(C,\Pi,\boldsymbol{f},\boldsymbol{g})$. An implicit differentiation approach is thus introduced in [20]. It implicitly differentiates the associated marginal constraints to derive a large linear system, which is solved by the conjugate gradient solver in lineax [26]. It applies to different regularized OT problems [10, 35, 36, 11, 6].
@@
For computing the Hessian of the OT distances, up to our knowledge, there is no direct analytical expression before the current work. There are two approaches suggested by OTT[9]. The first approach unrolls the Sinkhorn iterations and use the JAX in-build tools to handle autodiff via backward propagation and computational graph. The second approach implicitly differentiates the optimal solutions computed by OTT. The implicit differentiation approach involves differentiating the solution of an ill-conditioned linear system with the custom differentiation rules [9], hence, regularization techniques, such as preconditioning the marginal constraints [10] and the ridge regularization, have been introduced to try to resolve this issue. However, as we will demonstrate in Section 5, both approaches still encounter two major challenges: (i) memory shortages when the dataset is large, and (ii) numerical instability due to singularity and the ill-posed nature of the linear system, particularly when $\epsilon$ is small.
@@
The key in our derivation is the linear system for the optimal dual potentials, e.g., (26) or (34), emerged from the application of the implicit differentiation. It facilitates efficient computation of the gradient of the Sinkhorn distance as well as the Hessian of the EOT distance. Additionally, together with (29), it bridges implicit differentiation and Danskin’s theorem in the context of the EOT distance.
@@ 3.2 Analytical computation of the gradients
We first review the implicit differentiation approach to the gradient of the EOT distance $\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$ with respect to source data $\boldsymbol{Y}$ and re-derive the result of Danskin’s approach in (3.1) through the key observation. We further provide a novel numerical method to efficiently compute the derivative of Sinkhorn distance $\widetilde{\mathrm{OT}}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$.
Gradient of $\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})$. We first consider the gradient of the EOT distance with respect to the source data $\boldsymbol{y}_{k}$
$\frac{\partial\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{\partial\boldsymbol{y}_{k}}=\sum_{i=1}^{M}\mu_{i}\frac{\partial f_{i}^{*}}{\partial\boldsymbol{y}_{k}}+\sum_{j=1}^{N}\nu_{j}\frac{\partial g^{*}_{j}}{\partial\boldsymbol{y}_{k}}=\left[\begin{array}[]{ccccccccc}\boldsymbol{\mu}^{\top}&\boldsymbol{\nu}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\frac{\partial\boldsymbol{f}^{*}}{\partial\boldsymbol{y}_{k}}\\ \frac{\partial\boldsymbol{g}^{*}}{\partial\boldsymbol{y}_{k}}\end{array}\right],$ (16a)
where $\frac{\partial\boldsymbol{f}^{*}}{\partial\boldsymbol{y}_{k}}=\left(\frac{\partial f^{*}_{1}}{\partial\boldsymbol{y}_{k}},\ldots,\frac{\partial f^{*}_{M}}{\partial\boldsymbol{y}_{k}}\right)^{\top}\in\mathbb{R}^{M\times d}$, and similarly $\frac{\partial\boldsymbol{g}^{*}}{\partial\boldsymbol{y}_{k}}\in\mathbb{R}^{N\times d}$. To simplify the notation, we denote $\left(\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\right)_{iks}=\frac{\partial f^{*}_{i}}{\partial(\boldsymbol{y}_{k})_{s}}$ and $\left(\frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\right)_{jks}=\frac{\partial g^{*}_{j}}{\partial(\boldsymbol{y}_{k})_{s}}$, so $\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\in\mathbb{R}^{M\times M\times d}$ and $\frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\in\mathbb{R}^{N\times M\times d}$.
@@
$\frac{d\text{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{Y}}=\boldsymbol{\mu}^{\top}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}+\boldsymbol{\nu}^{\top}\frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}=\left[\begin{array}[]{ccccccccc}\boldsymbol{\mu}^{\top}&\boldsymbol{\nu}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\frac{d\boldsymbol{f}^{*}}{d\boldsymbol{Y}}\\ \frac{d\boldsymbol{g}^{*}}{d\boldsymbol{Y}}\end{array}\right]\in\mathbb{R}^{M\times d}.$ (16b)
@@
The derivative of EOT distance in (11) with respect to source data $\boldsymbol{y}_{k}$, as in (16a), is given by
$\frac{\partial\mathrm{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{\partial\boldsymbol{y}_{k}}=\sum_{j=1}^{N}\frac{\partial C_{kj}}{\partial{\boldsymbol{y}_{k}}}\Pi^{*}_{kj},\quad k=1,\dots,M.$ (17)
In vector form, the derivative of $\mathrm{OT}_{\epsilon}(C,\mu,\nu)$ with respect to whole source data $\boldsymbol{Y}$ is
$\frac{d\mathrm{OT}_{\epsilon}(C,\boldsymbol{\mu},\boldsymbol{\nu})}{d\boldsymbol{Y}}=\mathcal{B}\cdot\mathbbm{1}_{N},$ (18)
where $\mathcal{B}\in\mathbb{R}^{M\times d\times N}$ is a tensor with entries $\mathcal{B}_{ksj}=\frac{\partial C_{kj}}{\partial(\boldsymbol{y}_{k})_{s}}\Pi^{*}_{kj}$, and $\mathcal{B}\cdot\mathbbm{1}_{N}\in\mathbb{R}^{M\times d}$ is the dot product which is the summation of the third index such that the $k$-th column is $(\mathcal{B}\cdot\mathbbm{1}_{N})_{k}=\sum_{j}\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\Pi^{*}_{kj}$ for $1\leq k\leq M$.
@@
The main task is to find $\frac{\partial\boldsymbol{f}^{*}}{\partial\boldsymbol{y}_{k}}$ and $\frac{\partial\boldsymbol{g}^{*}}{\partial\boldsymbol{y}_{k}}$ in (16a) using (7) and the marginal probability conditions. The partial derivative of entries of optimal coupling matrix $\Pi^{*}$ with respect to $\boldsymbol{y}_{k}$ is
$\frac{\partial\Pi^{*}_{ij}}{\partial\boldsymbol{y}_{k}}$ $=\frac{\Pi^{*}_{ij}}{\epsilon}\left(-\frac{\partial C_{ij}}{\partial\boldsymbol{y}_{k}}+\frac{\partial f_{i}^{*}}{\partial\boldsymbol{y}_{k}}+\frac{\partial g_{j}^{*}}{\partial\boldsymbol{y}_{k}}\right)=\frac{\Pi^{*}_{ij}}{\epsilon}\left(-\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\delta_{ik}+\frac{\partial f_{i}^{*}}{\partial\boldsymbol{y}_{k}}+\frac{\partial g_{j}^{*}}{\partial\boldsymbol{y}_{k}}\right),$ (19)
where $\delta_{ik}$ is a Kronecker delta function. We observe that with the marginal probability conditions $\sum_{j=1}^{N}\Pi^{*}_{ij}=\mu_{i}$ and $\sum_{i=1}^{M}\Pi^{*}_{ij}=\nu_{j}$, and by taking the partial derivative $\frac{\partial}{\partial\boldsymbol{y}_{k}}$ on both sides of these marginal constraints, we can get
$0=\sum_{j=1}^{N}\frac{\partial\Pi^{*}_{ij}}{\partial\boldsymbol{y}_{k}}=\frac{\mu_{i}}{\epsilon}\frac{\partial f_{i}^{*}}{\partial\boldsymbol{y}_{k}}-\frac{1}{\epsilon}\left[\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\delta_{ik}-\frac{\partial g^{*}_{j}}{\partial\boldsymbol{y}_{k}}\right)\Pi^{*}_{ij}\right],$
$0=\sum_{i=1}^{M}\frac{\partial\Pi^{*}_{ij}}{\partial{\boldsymbol{y}_{k}}}=\frac{\nu_{j}}{\epsilon}\frac{\partial g_{j}^{*}}{\partial\boldsymbol{y}_{k}}-\frac{1}{\epsilon}\left[\sum_{i=1}^{M}\left(\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\delta_{ik}-\frac{\partial f^{*}_{i}}{\partial\boldsymbol{y}_{k}}\right)\Pi^{*}_{ij}\right].$
@@
$\mu_{i}\frac{\partial f_{i}^{*}}{\partial\boldsymbol{y}_{k}}+\sum_{j=1}^{N}\frac{\partial g^{*}_{j}}{\partial\boldsymbol{y}_{k}}\Pi^{*}_{ij}=\sum_{j=1}^{N}\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\Pi^{*}_{ij}\delta_{ik}\in\mathbb{R}^{d},\quad\forall 1\leq i\leq M,$
$\sum_{i=1}^{M}\frac{\partial f^{*}_{i}}{\partial\boldsymbol{y}_{k}}\Pi^{*}_{ij}+\nu_{j}\frac{\partial g_{j}^{*}}{\partial\boldsymbol{y}_{k}}=\frac{\partial C_{kj}}{\partial\boldsymbol{y}_{k}}\Pi^{*}_{kj}\in\mathbb{R}^{d},\quad\forall 1\leq j\leq N.$
With $H(\Pi)$ defined in (2), the above linear system can be written in matrix form,
@@
$\underbrace{\left[\begin{array}[]{ccccccccc}\text{diag}(\boldsymbol{\mu})&\Pi^{*}\\ (\Pi^{*})^{\top}&\text{diag}(\boldsymbol{\nu})\end{array}\right]}_{=H(\Pi^{*})}\left[\begin{array}[]{ccccccccc}\frac{\partial\boldsymbol{f}^{*}}{\partial\boldsymbol{y}_{k}}\\ \frac{\partial\boldsymbol{g}^{*}}{\partial\boldsymbol{y}_{k}}\end{array}\right]=\left[\begin{array}[]{ccccccccc}\mathbf{e}_{k}(B_{k}\mathbbm{1}_{N})^{\top}\\ B_{k}^{\top}\end{array}\right],$ (26)

### v2

@@ 3 Differentiation of Loss Functions
$\frac{\partial\text{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{\partial\theta_{i}}=\sum_{k}\frac{\partial{\mathbf{y}}_{k}}{\partial\theta_{i}}\frac{\partial\text{OT}_{\epsilon}(\bm{C}_{\theta},\bm{\mu},\bm{\nu})}{\partial{\mathbf{y}}_{k}},$ (12)
$\frac{\partial^{2}\mathrm{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{\partial\theta_{i}\partial\theta_{j}}=\sum_{s,k}\frac{\partial{\mathbf{y}}_{s}}{\partial\theta_{i}}\frac{\partial^{2}\mathrm{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{\partial{\mathbf{y}}_{s}\partial{\mathbf{y}}_{k}}\left(\frac{\partial{\mathbf{y}}_{k}}{\partial\theta_{j}}\right)^{\top}+\sum_{k}\frac{\partial^{2}{\mathbf{y}}_{k}}{\partial\theta_{i}\partial\theta_{j}}\frac{\partial\mathrm{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{\partial{\mathbf{y}}_{k}}.$
The key step is to find the explicit expression for the first and second derivatives with respect to the source data $\frac{\partial\text{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{\partial{\mathbf{y}}_{k}}\in\mathbb{R}^{d}$, $\frac{\partial^{2}\mathrm{OT}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})}{\partial{\mathbf{y}}_{s}\partial{\mathbf{y}}_{k}}\in\mathbb{R}^{d\times d}$. To provide a full picture, the gradients of EOT distance and Sinkhorn distance with respective to the measures $\bm{\mu}$ or $\bm{\nu}$ have been studied in [22, 26].
@@ 3.1 Previous methods computing first- and second-order derivatives
There are two main approaches to the analytical expression of the first-order derivatives of regularized OT distances: the Danskin approach and the implicit differentiation approach. The Danskin approach computes the gradient of EOT, based on applying the Danskin’s theorem to the dual function [5, 10, 17, 16].
Theorem 1 (Danskin’s theorem).
Suppose $\phi(\mathbf{x},\mathbf{z})$ is a continuous function of two arguments, $\phi:\mathbb{R}^{n}\times Z\rightarrow\mathbb{R}$, where $Z\subset\mathbb{R}^{m}$ is a compact set and define $f(\mathbf{x})=\max_{\mathbf{z}\in Z}\phi(\mathbf{x},\mathbf{z})$ and the set of maximizing points $\mathbf{z}^{*}(\mathbf{x})=\underset{\mathbf{z}\in Z}{\operatorname{arg}\operatorname{max}}\;\phi(\mathbf{x},\mathbf{z})$. If $\mathbf{z}^{*}(\mathbf{x})$ consists of a single element, the gradient of $f(\mathbf{x})$ is given by $\frac{\partial f}{\partial\mathbf{x}}=\frac{\partial\phi(\mathbf{x},\mathbf{z}^{*}(\mathbf{x}))}{\partial\mathbf{x}}$.
Recall the outer problem $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})=\max_{\mathbf{f},\mathbf{g}}L(\mathbf{C},\bm{\Pi}^{*}(\mathbf{f},\mathbf{g}),\mathbf{f},\mathbf{g})$,
$\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})=\max_{\mathbf{f},\mathbf{g}}\underbrace{\sum_{i}f_{i}\mu_{i}+\sum_{j}g_{j}\nu_{j}-\sum_{ij}\epsilon\mu_{i}\nu_{j}\exp\left(\frac{-C_{ij}+f_{i}+g_{j}}{\epsilon}\right)+\epsilon}_{\phi(\mathbf{Y},\mathbf{f},\mathbf{g})}.$
If $(\mathbf{f}^{*},\mathbf{g}^{*})$ is a single element, applying Danskin’s theorem to $\phi(\mathbf{Y},\mathbf{f},\mathbf{g})$, we obtain
$\frac{\partial\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{\partial\mathbf{y}_{k}}$ $=\sum_{ij}\mu_{i}\nu_{j}\exp\left(\frac{-C_{ij}+f^{*}_{i}+g^{*}_{j}}{\epsilon}\right)\frac{\partial C_{ij}}{\partial\mathbf{y}_{k}}=\sum_{j}\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\Pi^{*}_{kj}.$ (13)
@@
The last equality is due to $\frac{\partial C_{ij}}{\partial\mathbf{y}_{k}}=\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\delta_{ki}$. Once $\bm{\Pi}^{*}$ is obtained from the Sinkhorn algorithm, the first-order derivative with respect to the source dataset $\mathbf{Y}$ immediately follows without additional computational cost. Unfortunately Danskin’s approach doesn’t work on computing the gradient of the Sinkhorn distances $\widetilde{\text{OT}}_{\epsilon}(\mathbf{C}_{\theta},\bm{\mu},\bm{\nu})$ because it is not of the form $\max_{\mathbf{f},\mathbf{g}}\phi(\mathbf{C},\mathbf{f},\mathbf{g})$. An implicit differentiation approach is thus introduced in [22]. It implicitly differentiates the associated marginal constraints to derive a large linear system, which is solved by the conjugate gradient solver in lineax [28]. It applies to different regularized OT problems [12, 37, 38, 13, 7].
@@
For computing the Hessian of the OT distances, up to our knowledge, there is no direct analytical expression before the current work. Similarly, Danskin’s approach doesn’t work on computing the Hessian of EOT distance. There are two approaches suggested by OTT[11]. The first approach unrolls the Sinkhorn iterations and use the JAX in-build tools to handle autodiff via backward propagation and computational graph. The second approach implicitly differentiates the optimal solutions computed by OTT. The implicit differentiation approach involves differentiating the solution of an ill-conditioned linear system with the custom differentiation rules [11], hence, regularization techniques, such as preconditioning the marginal constraints [12, 39] and the ridge regularization, have been introduced to try to resolve this issue. However, as we will demonstrate in Section 5, both approaches still encounter two major challenges: (i) memory shortages when the dataset is large, and (ii) numerical instability due to singularity and the ill-posed nature of the linear system, particularly when $\epsilon$ is small.
@@
The key in our derivation is the linear system for the optimal dual potentials, e.g., (24) or (32), which arises from applying implicit differentiation. It facilitates efficient computation of the gradient of the Sinkhorn distance as well as the Hessian of the EOT distance. Additionally, together with (27), it bridges implicit differentiation and Danskin’s theorem in the context of the EOT distance.
@@ 3.2 Analytical computation of the gradients
We first review the implicit differentiation approach to the gradient of the EOT distance $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$ with respect to source data $\mathbf{Y}$ and re-derive the result of Danskin’s approach in (13) through the key observation. With the same technique, we further provide a novel numerical method to efficiently compute the derivative of Sinkhorn distance $\widetilde{\mathrm{OT}}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$.
Gradient of $\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$. We first consider the gradient of the EOT distance with respect to the source data $\mathbf{y}_{k}$
$\frac{\partial\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{\partial\mathbf{y}_{k}}=\sum_{i=1}^{M}\mu_{i}\frac{\partial f_{i}^{*}}{\partial\mathbf{y}_{k}}+\sum_{j=1}^{N}\nu_{j}\frac{\partial g^{*}_{j}}{\partial\mathbf{y}_{k}}=\left[\begin{array}[]{ccccccccc}\bm{\mu}^{\top}&\bm{\nu}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\frac{\partial\mathbf{f}^{*}}{\partial\mathbf{y}_{k}}\\ \frac{\partial\mathbf{g}^{*}}{\partial\mathbf{y}_{k}}\end{array}\right],$ (14a)
where $\frac{\partial\mathbf{f}^{*}}{\partial\mathbf{y}_{k}}=\left(\frac{\partial f^{*}_{1}}{\partial\mathbf{y}_{k}},\ldots,\frac{\partial f^{*}_{M}}{\partial\mathbf{y}_{k}}\right)^{\top}\in\mathbb{R}^{M\times d}$, and similarly $\frac{\partial\mathbf{g}^{*}}{\partial\mathbf{y}_{k}}\in\mathbb{R}^{N\times d}$. To simplify the notation, we denote $\left(\frac{d\mathbf{f}^{*}}{d\mathbf{Y}}\right)_{iks}=\frac{\partial f^{*}_{i}}{\partial(\mathbf{y}_{k})_{s}}$ and $\left(\frac{d\mathbf{g}^{*}}{d\mathbf{Y}}\right)_{jks}=\frac{\partial g^{*}_{j}}{\partial(\mathbf{y}_{k})_{s}}$, so $\frac{d\mathbf{f}^{*}}{d\mathbf{Y}}\in\mathbb{R}^{M\times M\times d}$ and $\frac{d\mathbf{g}^{*}}{d\mathbf{Y}}\in\mathbb{R}^{N\times M\times d}$.
@@
$\frac{d\text{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{d\mathbf{Y}}=\bm{\mu}^{\top}\frac{d\mathbf{f}^{*}}{d\mathbf{Y}}+\bm{\nu}^{\top}\frac{d\mathbf{g}^{*}}{d\mathbf{Y}}=\left[\begin{array}[]{ccccccccc}\bm{\mu}^{\top}&\bm{\nu}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\frac{d\mathbf{f}^{*}}{d\mathbf{Y}}\\ \frac{d\mathbf{g}^{*}}{d\mathbf{Y}}\end{array}\right]\in\mathbb{R}^{M\times d}.$ (14b)
@@
The derivative of EOT distance in (10) with respect to source data $\mathbf{y}_{k}$, as in (14a), is given by
$\frac{\partial\mathrm{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{\partial\mathbf{y}_{k}}=\sum_{j=1}^{N}\frac{\partial C_{kj}}{\partial{\mathbf{y}_{k}}}\Pi^{*}_{kj},\quad k=1,\dots,M.$ (15)
In vector form, the derivative of $\mathrm{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})$ with respect to whole source data $\mathbf{Y}$ is
$\frac{d\mathrm{OT}_{\epsilon}(\mathbf{C},\bm{\mu},\bm{\nu})}{d\mathbf{Y}}=\mathcal{B}\cdot\mathbb{1}_{N},$ (16)
where $\mathcal{B}\in\mathbb{R}^{M\times d\times N}$ is a tensor with entries $\mathcal{B}_{ksj}=\frac{\partial C_{kj}}{\partial(\mathbf{y}_{k})_{s}}\Pi^{*}_{kj}$, and $\mathcal{B}\cdot\mathbb{1}_{N}\in\mathbb{R}^{M\times d}$ is the dot product which is the summation of the third index such that the $k$-th column is $(\mathcal{B}\cdot\mathbb{1}_{N})_{k}=\sum_{j}\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\Pi^{*}_{kj}$ for $1\leq k\leq M$.
@@
The main task is to find $\frac{\partial\mathbf{f}^{*}}{\partial\mathbf{y}_{k}}$ and $\frac{\partial\mathbf{g}^{*}}{\partial\mathbf{y}_{k}}$ in (14a) and the marginal probability conditions. The partial derivative of entries of optimal coupling matrix $\bm{\Pi}^{*}$ with respect to $\mathbf{y}_{k}$ is
$\frac{\partial\Pi^{*}_{ij}}{\partial\mathbf{y}_{k}}$ $=\frac{\Pi^{*}_{ij}}{\epsilon}\left(-\frac{\partial C_{ij}}{\partial\mathbf{y}_{k}}+\frac{\partial f_{i}^{*}}{\partial\mathbf{y}_{k}}+\frac{\partial g_{j}^{*}}{\partial\mathbf{y}_{k}}\right)=\frac{\Pi^{*}_{ij}}{\epsilon}\left(-\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\delta_{ik}+\frac{\partial f_{i}^{*}}{\partial\mathbf{y}_{k}}+\frac{\partial g_{j}^{*}}{\partial\mathbf{y}_{k}}\right),$ (17)
where $\delta_{ik}$ is a Kronecker delta function. We observe that with the marginal probability conditions $\sum_{j=1}^{N}\Pi^{*}_{ij}=\mu_{i}$ and $\sum_{i=1}^{M}\Pi^{*}_{ij}=\nu_{j}$, and by taking the partial derivative $\frac{\partial}{\partial\mathbf{y}_{k}}$ on both sides of these marginal constraints, we can get
$0=\sum_{j=1}^{N}\frac{\partial\Pi^{*}_{ij}}{\partial\mathbf{y}_{k}}=\frac{\mu_{i}}{\epsilon}\frac{\partial f_{i}^{*}}{\partial\mathbf{y}_{k}}-\frac{1}{\epsilon}\left[\sum_{j=1}^{N}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\delta_{ik}-\frac{\partial g^{*}_{j}}{\partial\mathbf{y}_{k}}\right)\Pi^{*}_{ij}\right],$
$0=\sum_{i=1}^{M}\frac{\partial\Pi^{*}_{ij}}{\partial{\mathbf{y}_{k}}}=\frac{\nu_{j}}{\epsilon}\frac{\partial g_{j}^{*}}{\partial\mathbf{y}_{k}}-\frac{1}{\epsilon}\left[\sum_{i=1}^{M}\left(\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\delta_{ik}-\frac{\partial f^{*}_{i}}{\partial\mathbf{y}_{k}}\right)\Pi^{*}_{ij}\right].$
@@
$\mu_{i}\frac{\partial f_{i}^{*}}{\partial\mathbf{y}_{k}}+\sum_{j=1}^{N}\frac{\partial g^{*}_{j}}{\partial\mathbf{y}_{k}}\Pi^{*}_{ij}=\sum_{j=1}^{N}\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\Pi^{*}_{ij}\delta_{ik}\in\mathbb{R}^{d},\quad\forall 1\leq i\leq M,$
$\sum_{i=1}^{M}\frac{\partial f^{*}_{i}}{\partial\mathbf{y}_{k}}\Pi^{*}_{ij}+\nu_{j}\frac{\partial g_{j}^{*}}{\partial\mathbf{y}_{k}}=\frac{\partial C_{kj}}{\partial\mathbf{y}_{k}}\Pi^{*}_{kj}\in\mathbb{R}^{d},\quad\forall 1\leq j\leq N.$
The above linear system can be written in matrix form,
@@
$\left[\begin{array}[]{ccccccccc}\text{diag}(\bm{\mu})&\bm{\Pi}^{*}\\ (\bm{\Pi}^{*})^{\top}&\text{diag}(\bm{\nu})\end{array}\right]\left[\begin{array}[]{ccccccccc}\frac{\partial\bm{f}^{*}}{\partial\mathbf{y}_{k}}\\ \frac{\partial\bm{g}^{*}}{\partial\mathbf{y}_{k}}\end{array}\right]=\left[\begin{array}[]{ccccccccc}\mathbf{e}_{k}(\mathbf{B}_{k}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{k}^{\top}\end{array}\right],$ (24)
where the left-hand-side matrix is $\mathbf{H}^{*}$ defined in (2), $\mathbf{e}_{k}$ is the $k$-th standard column basis of $\mathbb{R}^{M}$ and the matrix $\mathbf{B}_{k}\in\mathbb{R}^{d\times N}$ is $(\mathbf{B}_{k})_{sj}=\frac{\partial C_{kj}}{\partial(\mathbf{y}_{k})_{s}}\Pi^{*}_{kj}$. Notice that $\left[\begin{array}[]{ccccccccc}\mathbb{1}_{M}^{\top}&-\mathbb{1}_{N}^{\top}\end{array}\right]\left[\begin{array}[]{ccccccccc}\mathbf{e}_{k}(\mathbf{B}_{k}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{k}^{\top}\end{array}\right]=\mathbb{0}_{d}$, and $\text{span}\left(\left[\begin{array}[]{ccccccccc}\mathbb{1}_{M}\\ -\mathbb{1}_{N}\end{array}\right]\right)$ is the kernel of $\mathbf{H}^{*}$ as proved in Theorem 6, so $\left[\begin{array}[]{ccccccccc}\mathbf{e}_{k}(\mathbf{B}_{k}\mathbb{1}_{N})^{\top}\\ \mathbf{B}_{k}^{\top}\end{array}\right]$ is in the span of $\mathbf{H}^{*}$, therefore the linear system above always has a solution.